# Investigating Feature Attribution for 5G Network Intrusion Detection

### v1

@@ Abstract
With the rise of fifth-generation (5G) networks in critical applications, it is urgent to move from detection of malicious activity to systems capable of providing a reliable verdict suitable for mitigation. In this regard, understanding and interpreting machine learning (ML) models’ security alerts is crucial for enabling actionable incident response orchestration. Explainable Artificial Intelligence (XAI) techniques are expected to enhance trust by providing insights into why alerts are raised. A dominant approach statistically associates feature sets that can be correlated to a given alert. This paper starts by questioning whether such attribution is relevant for future generation communication systems, and investigates its merits in comparison with an approach based on logical explanations. We extensively study two methods, SHAP and VoTE-XAI, by analyzing their interpretations of alerts generated by an XGBoost model in three different use cases with several 5G communication attacks. We identify three metrics for assessing explanations: sparsity, how concise they are; stability, how consistent they are across samples from the same attack type; and efficiency, how fast an explanation is generated. As an example, in a 5G network with 92 features, 6 were deemed important by VoTE-XAI for a Denial of Service (DoS) variant, ICMPFlood, while SHAP identified over 20. More importantly, we found a significant divergence between features selected by SHAP and VoTE-XAI. However, none of the top-ranked features selected by SHAP were missed by VoTE-XAI. When it comes to efficiency of providing interpretations, we found that VoTE-XAI is significantly more responsive, e.g. it provides a single explanation in under 0.002 seconds, in a high-dimensional setting (478 features).
@@ 1 Introduction
As in other domains, applying ML to security in 5G settings introduces a range of technical and practical challenges, and a gap remains between theoretical research and practical deployment [7]. The opaque, “black-box” nature of many ML models presents a significant obstacle in environments defined as high-risk under the European Union’s Artificial Intelligence (AI) Act [^1] https://digital-strategy.ec.europa.eu/en/policies/regulatory-framework-ai, including financial services, healthcare systems, and government infrastructures, many of which may be connected through wireless mobile interfaces.
@@
In this regard, tree ensembles offer a practical alternative to complex, black-box models due to their simplicity and clarity in structuring decision paths for classification and regression [13]. According to Chander et al. [14], decision trees can be employed when the features present in a dataset are dependent in nature, as in tabular data, which is the input type to network intrusion detection systems. Moreover, tree ensembles have demonstrated superior predictive performance over neural networks in several real-world, tabular-data applications [15].
@@
Note that the current paper does not aim to advance the state of intrusion detection in 5G networks, but assumes that further research can overcome some of the pitfalls of AI-based techniques for identifying attacks and anomalies [7].
Aiming to address further challenges, this study will provide a deeper understanding of the merits and limitations of explainability in the context of security monitoring within 5G networks.
@@
Several techniques are being explored for explainability, including Local Interpretable Model-Agnostic Explanations (LIME) [18], Anchors [19], Counterfactual [20], Integrated Gradient [21], and MaxSAT [22]. Multiple surveys cover the applicability of these methods in various domains and ML models [23, 24, 25, 26]. The large majority of the papers referred to in these surveys target classification tasks on images or text data. Since our aim is not to conduct a survey, we have selected representative statistical and logic-based approaches, suitable for tabular data. In particular, we chose methods based on feature attribution. These methods are particularly effective in models where understanding the influence of specific inputs is crucial [24].
@@
Existing statistical techniques, notably SHapley Additive exPlanations (SHAP) [27], have become a de facto standard for model interpretability by quantifying the contribution of each feature present in a prediction. However, recent studies [28, 29, 30] highlight potential shortcomings in the theoretical foundation of statistical methods. We are interested in understanding whether SHAP’s reliance on approximate Shapley values can yield misleading explanations, particularly by failing to capture meaningful relationships between input samples and features appearing in alerts. In cybersecurity, such inaccuracies are critical, as flawed explanations can lead to missed causes, or delayed, ineffective, and even harmful responses.
@@
Logic-based explainability methods  [28], and in particular those operating on tree ensembles, such as VoTE-XAI [31], have been proposed to provide minimal, provably correct explanations. Since logic-based explanations are known to have polynomial-time complexity in the size of the model [32, 33], it is interesting to assess whether they are useful in a high-dimensional domain like 5G networking. Given that network outages have been shown to last over a day [^2] https://www.vodafone.pt/en/press-releases/2022/2/cyberattack-on-vodafone-portugal.html, we believe that studying the capabilities of such methods towards rapid root-cause identification and service restoration is worth attention. Yet, the potential of logic-based explainability in live incident response remains largely unexplored.
The contribution of this paper is two-fold.
@@
    We thoroughly analyze the merits of logic-based and statistical interpretations of alerts based on feature attribution, with data from three different 5G testbeds, focusing on conciseness and stability across multiple samples. In particular, we:
    • (a)
      show the conciseness of the minimal explanations by VoTE-XAI even in high-dimensional settings,
    • (b)
      identify instability in SHAP’s feature attribution compared to VoTE-XAI,
    • (c)
      illustrate divergence among the methods in terms of attack-specific features consistently captured by VoTE-XAI but overlooked by SHAP.
@@
    Trade-off analysis between explainability and computational efficiency in the 5G context. Specifically, we:
    • (a)
      compare runtime performance, showing VoTE-XAI yields faster single (minimal) explanation outputs,
    • (b)
      quantify the cost of generating all minimal explanations, potentially useful in a forensic analysis.
@@
The remainder of the paper is organized as follows: Section 2 provides an overview of the feature attribution methods applied in the paper; Section 3 shows the proposed workflow, while Section 4 provides details regarding the use cases. The evaluation criteria are shown in Section 5, and the results are detailed in Section 6. The outcomes, threats to validity, and future directions are discussed in Section 7. Section 8 reviews the related work. Finally, Section 9 ends the paper with closing remarks.
@@ 2 Background
XAI refers to a suite of techniques that aim to enhance transparency and improve user trust in ML models by adding accountability to their decision-making processes and providing insights into how predictions are generated. The need for explainability was recognized long before the term XAI was introduced in 2016, with early works already highlighting challenges in making model reasoning accessible to end users [34].
Although there is no general agreement on how explainability shall be understood, the methods can be classified along multiple dimensions, including the scope of explanation, the stage of application, and the dependency on model architecture [35, 25]. For instance, global explanations offer insights into the overall decision-making behavior of an ML model, while local explanations focus on individual outputs. Another classification considers whether a technique depends on a model’s internal structure. Model-agnostic methods explain predictions using only inputs and outputs. Model-specific methods, in contrast, leverage the internal mechanisms of particular ML architectures.
In this research, we apply two explainability techniques for local explanations: SHAP, which estimates feature importance using statistical approaches, and for which we rely on the TreeExplainer optimization [^3] https://shap.readthedocs.io/en/latest/generated/shap.TreeExplainer.html, and VoTE-XAI, a logic-based explanation technique designed for tree ensembles.
2.1 SHAP
SHAP is an explainability technique grounded in cooperative game theory. It explains individual predictions by assigning importance scores, known as Shapley values, to input features. These scores represent the contribution of each feature to the model’s output.
The Shapley value $\phi_{i}$, for a feature $i$ is defined as shown in Equation 1:
@@ 2.1 SHAP
Where $F$ is the set of all features, $S\subseteq F\setminus\{i\}$ is a subset of features excluding $i$, $f_{S\cup\{i\}}$ is the model trained with features in $S\cup\{i\}$, while $f_{S}$ is the model trained with only features in $S$. Finally, $x_{S\cup\{i\}}$ and $x_{S}$ are the input values for the respective feature subsets.
SHAP is generally model-agnostic, as it can be applied to a wide range of ML models. However, optimized implementations exist for specific architectures, such as the TreeExplainer module for tree ensembles. This method is widely adopted due to its scalability and ability to provide both global and local explanations of model behavior.
2.2 VoTE-XAI
VoTE-XAI focuses on generating explanations for predictions made by tree ensemble models. The approach provides explanations that are provably correct and minimal with respect to a given prediction inferred from the model. Specifically, an explanation $E$ is valid for a prediction $f(c_{1},\dots,c_{n})\mapsto d$ if:

### v2

# Feature Attribution in 5G Intrusion Detection: A Statistical vs. Logic-Based Comparison

@@ Abstract
With the rise of fifth-generation (5G) networks in critical applications, it is urgent to move from detection of malicious activity to systems capable of providing a reliable verdict suitable for mitigation. In this regard, understanding and interpreting machine learning (ML) models’ security alerts is crucial for enabling actionable incident response orchestration. Explainable Artificial Intelligence (XAI) techniques are expected to enhance trust by providing insights into why alerts are raised. Under the umbrella of XAI, interpretability of outcomes is crucially dependent on understanding the influence of specific inputs, referred to as feature attribution. A dominant approach to feature attribution statistically associates feature sets that can be correlated to a given alert. This paper investigates its merits against the backdrop of criticism from recent literature, in comparison with feature attribution based on logic. We extensively study two methods, SHAP and VoTE-XAI, as representatives of each feature attribution approach by analyzing their interpretations of alerts generated by an XGBoost model across three 5G-relevant datasets (5G-NIDD, MSA, and PFCP) covering multiple attack scenarios. We identify three metrics for assessing explanations: sparsity, how concise they are; stability, how consistent they are across samples from the same attack type; and efficiency, how fast an explanation is generated. Our results reveal that logic-based attributions are consistently more sparse and stable across alerts. More importantly, we found a significant divergence between features selected by SHAP and VoTE-XAI. However, none of the top-ranked features selected by SHAP were missed by VoTE-XAI. Finally, we analyze the efficiency of both methods, discussing their suitability for real-time security monitoring even in high-dimensional 5G environments (478 features).
@@ 1 Introduction
As in other domains, applying ML to security in 5G settings introduces a range of technical and practical challenges, and a gap remains between theoretical research and practical deployment [5].
The opaque, “black-box” nature of many ML models presents a significant obstacle in environments defined as high-risk under the European Union’s Artificial Intelligence (AI) Act [^1] https://digital-strategy.ec.europa.eu/en/policies/regulatory-framework-ai, including financial services, healthcare systems, and government infrastructures, many of which may be connected through wireless mobile interfaces.
@@
In this regard, tree ensembles offer a practical alternative to complex, black-box models due to their simplicity and clarity in structuring decision paths for classification and regression [51]. According to Chander et al. [11], decision trees can be employed when the features present in a dataset are dependent on each other, as in tabular data, which is the input type to network intrusion detection systems. Moreover, tree ensembles have demonstrated superior predictive performance over neural networks in several real-world, tabular-data applications [52].
@@
This paper addresses the trustworthiness of alarms created based on tabular data within a security monitoring context for critical infrastructure. It aims to provide a deeper understanding of the merits and limitations of representative methods for interpretability within 5G networks. Note that the current paper does not aim to advance the state of intrusion detection in 5G networks, but assumes that further research can overcome some of the pitfalls of AI-based techniques for understanding attacks and anomalies [5].
@@
Several techniques are being explored under the umbrella of explainability, including SHapley Additive exPlanations (SHAP) [30], Local Interpretable Model-Agnostic Explanations (LIME) [45], Anchors [46], Counterfactual (CF) [59], Integrated Gradient (IG) [54], MaxSAT [26], and VoTE-XAI [56]. Multiple surveys cover the applicability of these methods in various domains and ML models [53, 41, 50, 34]. The large majority of the papers referred to in these surveys target classification tasks on images or text data. Since our aim is not to conduct a survey, we have selected representative statistical and logic-based approaches, suitable for tabular data. Furthermore, we have selected methods based on feature attribution. These are particularly effective in applications where understanding the influence of specific inputs is crucial [41].
Hence, we have excluded those methods whose interpretability goals do not align with feature attribution, such as CF (changes in feature values required to flip the model output) and Anchors (that generates decision rules). We have excluded IG since we restrict our study to non-differentiable ML models. While IG implementations for non-differentiable models exist [3], they are based on approximations, which might hinder the fairness of the comparison. We have excluded LIME since comparative analysis with SHAP reveals that the latter exhibits better quality according to the parameters defined by the authors [49]. Finally, we have excluded MaxSAT, since in comparison to VoTE-XAI against standard AI benchmarks, it has shown slower runtime behavior [56].
@@
Existing statistical techniques, notably SHAP, have become a de facto standard for model interpretability by quantifying the contribution of each feature present in a prediction. However, recent studies [33, 25, 31] highlight potential shortcomings in the theoretical foundation of statistical methods. We are interested in understanding whether SHAP’s reliance on approximate Shapley values can impact the capture of meaningful relationships between input samples and features appearing in alerts. In cybersecurity, inaccuracies are critical, as flawed attributions may lead to missed causes, or delayed, ineffective, and even harmful responses.
@@
Logic-based methods  [33], and in particular those operating on tree ensembles, such as VoTE-XAI, have been proposed to provide minimal, provably correct explanations. Since logic-based explanations are known to have polynomial-time complexity in the size of the model [8, 32], it is interesting to assess whether they are useful in a high-dimensional domain like 5G networking. Given that network outages have been shown to last over a day [^2] https://www.vodafone.pt/en/press-releases/2022/2/cyberattack-on-vodafone-portugal.html, we believe that studying the capabilities of such methods towards rapid root-cause identification and service restoration is worth attention. Yet, the potential of logic-based approaches in live incident response remains largely unexplored.
The objective of this study is to evaluate the merits of the selected feature attribution methods in 5G security monitoring. To enable comparison, we focus on three metrics (described in Section 6) that reflect operational requirements. Sparsity [60] captures the conciseness of attribution, important for interpretability in high-dimensional settings and further mitigation actions. Stability [1] measures the consistency of attribution across alerts of the same attack type, useful to assess whether similar inputs yield similar outputs. Efficiency [60] reflects the computational cost of generating attributions and their suitability for real-time security workflows.
The contributions of the paper are as follows:
@@
    We thoroughly analyze the merits of logic-based and statistical interpretations of alerts in a 5G setting based on feature attribution. The evaluation is carried out with data from three different testbeds, focusing on sparsity and stability across multiple samples.
@@
    We identify the degree of alignment between statistical and logic-based attributions in terms of attack-specific features.
  3. 3.
    We measure efficiency under two complementary perspectives: a single explanation in real-time contexts, and generating all logic-based explanations, potentially relevant in forensic analysis.
@@
The remainder of the paper is organized as follows: Section 2 reviews the related work; Section 3 provides an overview of XAI and the feature attribution methods studied in the paper; Section 4 shows the proposed method, while Section 5 provides details regarding the use cases. The evaluation criteria are shown in Section 6, and the results are detailed in Section 7. The outcomes, threats to validity, and future directions are discussed in Section 8. Finally, Section 9 ends the paper with closing remarks.
2 Related Work
A large body of existing studies on XAI focuses on statistical feature attribution, particularly SHAP and LIME, with limited exploration of logic-based methods in cybersecurity. In this section, we summarize relevant research and position our study within this landscape. Note that in the following, we use the term ”XAI” to refer to a plethora of methods for interpretability that are not based on feature attribution. Similarly, we will use the word ”explanation” to refer to the output of XAI techniques.
2.1 Explainable AI for Intrusion Detection
Various approaches have been applied to intrusion detection systems (IDS) and security monitoring to enhance interpretability and improve trust in ML-based threat detection. These studies demonstrate the practical value of applying XAI to IDS. However, in these works, the XAI methods themselves are not the primary object of analysis.
Gaspar et al. [21] explore SHAP and LIME to interpret predictions from Multi-Layer Perceptron (MLP)-based IDSs. Their work aims to evaluate the impact of feature attribution on the accuracy of the classifier. A perturbation analysis was conducted by removing or replacing features based on their importance based on SHAP/LIME and observing changes in classification. The study found that, in certain scenarios, relying on the methods negatively impacted the model’s performance, suggesting potential inconsistencies in the attributions. While this work uncovers possible limitations of LIME and SHAP, it neither explicitly highlights nor addresses them. Complementarily, our work attempts to gain a deeper understanding of the consequences of statistical and logic-based attribution.
Gyawali et al. [22] combine SHAP, LIME, and CF to enhance interpretability in IoT security. By employing these techniques, the researchers seek to translate model predictions into actionable strategies for attack mitigation. While the results are encouraging, the resulting explanations are implicitly treated as ground truth. Before linking attribution to actions, a critical step would be assessing their accuracy with respect to the data and the context, which is the aim of the present work.
Kalutharage et al. [27] propose a neurosymbolic approach to IoT security, integrating autoencoder-based anomaly detection with SHAP explanations and domain knowledge encoded in a knowledge graph. Similarly, Choi et al. [12] develop an attack-specific feature analysis framework that leverages SHAP to optimize feature selection for unsupervised models and autoencoders. In both works, attribution is used to refine detection mechanisms or improve alert generation prior to or during model deployment. In contrast, our work focuses on interpreting alerts in operational settings, aiming to support post-detection analysis and decision-making.
2.2 Explainable AI and Model Performance
Beyond IDS, multiple works leverage XAI as a tool to improve or diagnose the underlying ML models.
Najibi et al. [39] apply SHAP and LIME to extract the most influential features from deep neural network classifications and use domain knowledge to identify features considered more resistant to evasion. The models are then retrained using only these selected features, resulting in degraded detection performance. Vo et al. [58] similarly use SHAP for dimensionality reduction across various ML models by retraining classifiers on top-ranked features, observing performance degradation in some cases and slight improvements in others. Both studies show that attribution-guided feature selection led to performance degradation in certain scenarios, but the explicit investigation of this behavior was not part of the scope. In contrast, our work aims at investigating feature attribution per se.
Drichel et al. [15] used white-box, gradient-based techniques from the iNNvestigate library [^3] https://github.com/albermax/innvestigate to analyze deep learning models trained to reveal botnet activity through Domain Generation Algorithm (DGA) detection. Their goal is to use XAI to uncover bias in the classifier, showing a degradation of the model’s performance once the bias is removed. While their findings highlight the value of such techniques for understanding classifier behavior, the study uses them as diagnostic tools for model analysis. We apply feature attribution methods with a different scope, assessing their merits for post-detection interpretation.
2.3 Explainable AI for 5G and Beyond
Recent works have explored the integration of XAI into 5G/6G architectures, motivated by operational constraints such as scalability, latency, and interpretability of network-level decisions.
Basaran et al. [7] present XAInomaly, an anomaly detection framework for 6G networks. They propose an extension of SHAP designed to operate within the O-RAN architecture. Their work focuses on attribution for intelligence orchestration and network operation, emphasizing real-time feasibility through runtime and resource usage assessment. Terra et al. [55] apply global and local techniques to analyze SLA violations in 5G network slicing, validating their results with causal inference tools. While a diverse set of XAI techniques is applied, the research does not explore logic-based approaches. In contrast to both works, our study explicitly compares statistical and logic-based feature attributions in a security monitoring setting rather than network control.
2.4 Studies on Feature Attribution
Several works have examined statistical methods, particularly SHAP and LIME, in networking and security settings. These studies raise concern on conceptual and practical shortcomings of statistical attributions, motivating the need for alternative explanation paradigms.
Fiandrino et al. [18] demonstrate that SHAP, when applied out-of-the-box, produces non-intuitive and costly attributions for deep reinforcement learning (DRL) agents in O-RAN systems, as it relies on intermediate representations rather than meaningful inputs. While they propose a graph-based alternative, the scope is limited to network control and does not extend to intrusion detection or broader security monitoring.
Huang and Marques-Silva [25] critically analyze SHAP’s shortcomings in security settings, demonstrating that its assumptions of feature independence can lead to misleading attributions in complex datasets. The authors argue for the development of more reliable methods that accurately reflect feature relevance, but proposing an alternative was not in the scope of their work. This highlights the necessity for alternative approaches that do not rely on statistical assumptions.
Marques-Silva [32] gave a tutorial in a summer school, highlighting that statistical methods like SHAP and LIME often yield inconsistent and logically unsound attributions using small-scale synthetic models. They suggest that abductive and contrastive attribution, computed via Boolean Satisfiability (SAT) and Maximum Satisfiability (MaxSAT), can ensure rigor and minimality in explanations.
Our work takes this theoretical observation further to evaluate logic-based feature attribution in a practical landscape, applies it to 5G security, and further studies the timeliness of the explanation computations. Compared to the above state of the art, we empirically evaluate SHAP and VoTE-XAI based on sparsity, stability, and efficiency in three 5G security use cases.
Finally, we position our work compared to the work introducing VoTE-XAI. Törnblom et al. [56] focus on the theoretical underpinnings of the method to derive the explanations in tree ensembles, the soundness and completeness proofs, and some experiments in order to compare the approach to MaxSAT-based approaches. In these experiments, VoTE-XAI showed superior performance. However, the experimental evaluation is limited to common AI benchmarks, with fewer features (ranging from 7 to 60 as the highest dimension). As far as we are aware, the present work is the first to apply logic-based attributions in intrusion detection with a realistic experimental evaluation.
Considering the wide range of works we have referred to above, we believe that the two representative approaches with their inherent differences, deserve a detailed analysis in realistic case studies on fair grounds. While SHAP and Vote-XAI approaches to attribution are different in nature, we have strived to perform comparative analysis of the usefulness of each on fair grounds.
@@ 3 Background
In this section, we provide an overview on XAI, followed by concepts of statistical and logic-based feature attribution, with a focus on the methods analyzed in our study.
3.1 Explainable AI Scope and Concepts
XAI is an umbrella term which refers to a suite of techniques that aim to enhance transparency, and improve user trust in ML models by adding accountability to their decision-making processes and providing insights into how predictions are generated. The need for explainability was recognized long before the term XAI was introduced in 2016, with early works already highlighting challenges in making model reasoning accessible to end users with precursors presented in a review by Främling [19].
Beyond transparency, XAI aims at other goals, including understanding model behavior, diagnosing errors and biases, enabling informed decision-making, and ensuring that explanations are meaningful for different target audiences. At the same time, open challenges exist, such as lack of rigorous evaluation criteria for explanation quality and the risk of misleading explanations when underlying assumptions are violated [6].
Although there is no general agreement on how explainability should be understood, methods to interpret ML models can be classified along multiple dimensions, including the scope of explanation, the stage of application, and the dependency on model architecture [6, 50]. For instance, global explanations offer insights into the overall decision-making behavior of an ML model, while local explanations focus on individual outputs. Techniques can also be classified based on when they are applied in the ML pipeline: post-hoc methods provide explanations after the model has been trained and deployed. In contrast, ante-hoc refers to building the ability to interpret into a given model. Another classification considers whether a technique depends on a model’s internal structure. Model-agnostic methods explain predictions using only inputs and outputs. Model-specific methods, in contrast, leverage the internal mechanisms of particular ML architectures.
Within the broad landscape of XAI, feature attribution methods [41] constitute a class of post-hoc methods, aiming to identify subsets of input features to justify each model prediction.
Table 1 summarizes the concepts of statistical and logic-based feature attribution, as further detailed in the following subsections.
Table 1: Conceptual comparison between statistical and logic-based feature attribution.
Aspect Statistical Attribution Logic-based Attribution
Interpretability
goal
Which features influence the
model output, and by how
much?
Which features and values justify
the model output?
Definition of explanation
Numerical feature importance
scores to estimate the
contribution of individual
features to the predictions
Minimal sets of feature–value
pairings sufficient to justify the
given prediction
Scope
Local explanations, can be
aggregated to derive global
trends.
Local, tied to individual model
predictions
Model
dependency
Model-agnostic, with
optimized variants for specific
architectures
Model-specific, requiring access
to the model internals
Output
Single attribution vector per
prediction
Single or multiple attributions per
prediction. Multiple attributions
can be compared to each other if
preferences (domain-based weights
for features) exist
3.2 Statistical Feature Attribution and SHAP
Statistical feature attribution methods form a prominent class of post-hoc techniques. Representative approaches include additive methods such as SHAP and LIME [30, 45]. Comprehensive surveys report that these methods are widely applied across domains and model classes, with a strong emphasis on classification tasks involving image and text data [53, 41, 50, 34].
The de facto standard, SHAP, is a technique grounded in cooperative game theory. It explains individual predictions by assigning importance scores, known as Shapley values, to input features. These scores represent the contribution of each feature to the model’s output.
The Shapley value $\phi_{i}$, for a feature $i$ is defined as shown in Equation 1 [30]:
@@ 3.2 Statistical Feature Attribution and SHAP
Where $F$ is the set of all features, $S\subseteq F\setminus\{i\}$ is a subset of features excluding $i$, $f_{S\cup\{i\}}$ is the model trained with features in $S\cup\{i\}$, while $f_{S}$ is the model trained with only features in $S$. Finally, $x_{S\cup\{i\}}$ and $x_{S}$ are the input values drawn from the respective feature subsets.
SHAP is generally model-agnostic, as it can be applied to a wide range of ML models. However, optimized implementations exist for specific architectures, such as the TreeExplainer [^4] https://shap.readthedocs.io/en/latest/generated/shap.TreeExplainer.html module for tree ensembles. This method is widely adopted due to its scalability and ability to provide both global and local attributions of model behavior.
SHAP attributions are typically interpreted through visualizations to convey feature contributions at different levels of granularity. Local attributions are commonly shown using force plots or waterfall plots, which illustrate how individual feature contributions combine to produce a specific output. Global attribution is often summarized using beeswarm plots, which aggregate local attributions across many instances to highlight overall feature importance and variability.
3.3 Logic-based Feature Attribution and VoTE-XAI
Logic-based (or formal) feature attribution methods aim at computing explanations that are logically sound with respect to a model’s prediction. These approaches typically rely on Boolean reasoning, satisfiability (SAT), or optimization variants such as MaxSAT [26][33] or abstract interpretations [56]. Within this category, attributions are defined in terms of feature-value pairs that are sufficient to entail a prediction, rather than estimated contribution scores.
VoTE-XAI [56] focuses on generating explanations for predictions made by tree ensemble models. The approach provides explanations that are provably correct and minimal with respect to a given prediction inferred from the model.
Specifically, an explanation $E$ is valid for a prediction $f(c_{1},\dots,c_{n})\mapsto d$ if: